Mrs. Smith is 2 years more than 3 times as old as her son. If the difference between their ages is 22 years, how old is Mrs. Smith?
will the answer be 32

Answer by stanbon(75887)

Mrs. Smith is 2 years more than 3 times as old as her son. If the difference between their ages is 22 years, how old is Mrs. Smith?
-----------------
Let Mrs. Smith's age be x yrs.
Then her son is "x-22" yrs old.
EQUATION:
x-2=3(x-22)
x-2=3x-66
2x=64
x=32 (Mrs. Smith's age)
x-22=10 (Her son's age)
-------
Cheers,
Stan H.